# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v2

@@ III Proposed Method
Let $\mathbf{D}_{0}\in\mathbb{R}^{H\times W}$ be the depth map scanned from an object. After back-projection, $\mathbf{D}_{0}$ yields a partial point cloud, denoted as $\mathbf{P}_{in}\in\mathbb{R}^{N\times 3}$. Denote by $\mathbf{P}_{com}\in\mathbb{R}^{M\times 3}$ the complete shape of $\mathbf{P}_{in}$, where $H$ and $W$ represent the height and width of the depth image, $N$ and $M$ are the numbers of points. Our goal is to find a mapping $f(\mathbf{P}_{in})=\mathbf{P}_{out}$ to make the predicted complete shape $\mathbf{P}_{out}$ as close as possible to $\mathbf{P}_{com}$ without $\mathbf{P}_{com}$ as supervision.
As depicted in Figure 1, our framework, namely MAL-SPC, first leverages a Pattern Retrieval Network to generate complete point clouds. Then, we render the complete predictions into a silhouette map and multi-view depth maps. The rendered depth maps are used to perform adversarial learning with the counterparts from the same categories. In what follows, we will provide a detailed explanation of our approach.
@@ III-A Pattern Retrieval Network
The Pattern Retrieval Network leverages objects’ region-level geometric similarity to infer the missing structures. As shown in Figure 2, consuming as input a partial point cloud $\mathbf{P}_{in}$, we first use a 3D encoder derived from PCN [15] to learn a global feature $\mathbf{o}$ that encapsulates the object’s shape information. Then we employ three layers of fully-connected layers to reconstruct a coarse yet complete point cloud $\mathbf{P}_{c}\in\mathbb{R}^{N\times 3}$. Next, we extract the position encodings and curvature encodings from the partial input and coarse output, respectively. The position encodings represent the positional relationship of local points, which are formulated as the average distance between each point and its nearest neighbors. Specifically, for each point $\mathbf{p}_{i}\in\mathbf{P}_{in}$, we first find out its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})=\{\mathbf{q}_{i}^{j}\}_{j=1}^{K}\subset\mathbf{P}_{in}$ and then compute the average Euclidean distance between $\mathbf{p}_{i}$ and $\mathcal{N}_{K}(\mathbf{p}_{i})$, i.e.,
@@
Figure 2: The flowchart of our Pattern Retrieval Network. “FC” and “MLP” denote the Fully-Connected layer and the Multi-Layer Perceptron layer. Given a partial point cloud $\mathbf{P}_{in}$, we first utilize an Auto-encoder to generate a coarse shape $\mathbf{P}_{c}$. Next, we separately extract position encodings and curvature encodings for both $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$, and then leverage a cross-attention-based module to identify and retrieve the most related encodings for each point in $\mathbf{P}_{c}$. Subsequently, $\mathbf{P}_{c}$ and the selected encodings are upsampled and fused to refine the shape.
The curvature encodings capture the curvature variations across the surfaces of the point cloud. We employ the normals of individual points to characterize surface curvature. For each point $\mathbf{p}_{i}$, we first compute the covariance matrix $\mathbf{C}_{i}$ based on its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})$ determined previously. The covariance matrix captures the local surface geometry and is defined as follows:
@@
To simulate the curvature variations of the surface, inspired by [67], we calculated the cosine similarity between the normal vectors of each point and its neighboring $K$ points.
@@
where $\epsilon$ is a constant to prevent the denominator from being 0.
@@
In this way, we obtain the position encoding $\mathbf{f}_{pos}^{in}\in\mathbb{R}^{N\times 1}$ and curvature encoding $\mathbf{f}_{cur}^{in}\in\mathbb{R}^{N\times 1}$ from the partial input. Similarly, we compute the position encoding $\mathbf{f}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and curvature encoding $\mathbf{f}_{cur}^{c}\in\mathbb{R}^{N\times 1}$ from the coarse shape using the same method. Following this, we introduce a cross-attention-based pattern retrieval module to capture correlated local patterns for coarse predictions from partial inputs. Specifically, for the position encoding, we utilize MLPs to project $\mathbf{f}_{pos}^{in}$ and $\mathbf{f}_{pos}^{c}$ to get queries $\mathbf{g}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and keys $\mathbf{h}_{pos}^{in}\in\mathbb{R}^{N\times 1}$, respectively. Then we employ the algorithm in cross-attention [51] to calculate the weight maps between $\mathbf{h}_{pos}^{in}$ and $\mathbf{g}_{pos}^{c}$, representing the correlation of the position encoding for each point in $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$:
Next, based on $\mathbf{W}_{pos}$, we retrieve the $L$ most correlated position encodings from $\mathbf{f}_{pos}^{in}$ for each point in $\mathbf{P}_{c}$, denoted as $\mathbf{e}_{pos}\in\mathbb{R}^{N\times L}$. Similarly,the related curvature encodings $\mathbf{e}_{cur}\in\mathbb{R}^{N\times L}$ for each point in $\mathbf{P}_{c}$ are computed in the same manner. Then we upsample $\mathbf{P}_{c}$ along with $\mathbf{e}_{pos}$, $\mathbf{e}_{cur}$, $\mathbf{f}_{pos}^{c}$, $\mathbf{f}_{cur}^{c}$, $\mathbf{o}$, and fuse them in a concatenation manner. Finally, we employ MLPs to infer the point-wised offsets for the upsampled $\mathbf{P}_{c}$ and add them to obtain the final outputs $\mathbf{P}_{out}$.
@@ III-B Multi-view Adversarial Learning
Objects of the same category share many similar structures, and therefore, their shape distributions are close. To make the generated complete point cloud as realistic and reasonable as possible, we develop an adversarial learning module to learn the category-specific shape knowledge from the real depth maps. Specifically, we first employ a silhouette renderer to render $\mathbf{P}_{out}$ and $\mathbf{P}_{c}$ into two silhouette maps $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ from the same viewpoint as $\mathbf{D}_{0}$. $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ are supervised by $\mathbf{D}_{0}$. Then we leverage a depth renderer to render $\mathbf{P}_{out}$ into depth maps from different viewpoints. These depth maps and those from the category-specific depth image bank are fed to a discriminator, which consists of five 2D convolutional layers, to discriminate whether they are generated or real samples. It is important to note that the depth image bank is composed of single-view maps, each corresponding one-to-one with the input partial shape, without including samples from additional viewpoints.
@@
Our differentiable rendering pipeline is derived from [70]. Given the intrinsic and extrinsic camera parameters, [70] first transform the point cloud from world to camera coordinates and then project it into the 2D image plane. During rasterization, each point is rendered as a circular splat with a predefined radius in screen space, where the opacity of the splat decreases away from its center. The value of each pixel in the silhouette maps is computed by blending the information from the K nearest points along the z-axis whose splat regions overlap that pixel. For depth maps, each pixel’s value is the smallest z-coordinate of the K nearest points.
@@
However, [70] assigns a fixed and identical radius to each point splat. In fact, the density of 2D points projected by a point cloud with a fixed point number from different viewpoints is different. For example, as shown in Figure 3, when there is a large foreground area of a rendered image, the density of the projected points is small (we call these low-density depth maps). In such cases, a small radius cannot cover all areas of the visible surface and thus many “hole pixels” will appear in the rendered depth map. These “hole pixels” are inaccurately represented by the depth values of the points that should have been occluded. In contrast, in scenarios with few foreground pixels, the density of points is large (we call these high-density depth maps). In this case, a small radius suffices to cover the entire visible surface, and thus, the rendered depth map is more realistic and smooth. Therefore, a fixed point radius is inadequate for rendering under different viewpoints. Intuitively, using a larger point radius might produce smoother depth maps in both high point density and low point density scenarios. However, in this way, the projected points cover a larger area on the 2D plane, potentially extending the edges of the foreground pixels beyond the original boundaries of the object.
@@
To this end, we introduce a density-aware radius estimation (DARE) algorithm to adaptively determine a suitable radius for points from each viewpoint. First, we predefine an initial radius for the generated points and then render them into a coarse depth map. Note that the predefined point radius should cover the projected area. We then calculate the number of foreground pixels on this depth map, denoted as $A$, and subsequently, the density of the projected points on the depth map is defined as
@@ III-C Loss Functions
Furthermore, we use a rendering loss to supervise the predicted point cloud in 2D space. First, we binarize the depth image $\mathbf{D}_{0}$ to obtain the 2D silhouette ground truth $\mathbf{S}_{0}$ from the input viewpoint. Subsequently, we compute the Mean Squared Error (MSE) loss between $\mathbf{S}_{0}$ and $\mathbf{S}_{out}$. Note that due to the sparsity of the coarse point cloud $\mathbf{P}_{c}$, the foreground pixels of $\mathbf{S}_{c}$ are also sparse and discrete, whereas the foreground of $\mathbf{S}_{0}$ is continuous and dense. To solve this problem, we propose a Masked MSE loss, which only calculates the MSE loss on the foreground pixels of the rendered image. This trick confines the projected points of the rendered image to be accurately positioned.
$\begin{split}L_{rend}&=\frac{1}{H\cdot W}\sum_{i=0}^{H}\sum_{j=0}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{out}(i,j))^{2}\\ &+\frac{1}{H\cdot W}\sum_{i=0}^{H}\sum_{j=0}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{c}(i,j))^{2}\cdot\mathbf{B}(i,j),\end{split}$ (10)
@@
To achieve a more uniform point cloud, following [30], we incorporate a density loss $L_{dens}$ to constrain the local density of the point cloud to be as similar as possible.
@@
Methods Supervised Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
FoldingNet [41] full 6.41 1.86 9.13 3.19 5.72 12.72 5.42 7.29 6.00
PCN [15] full 5.42 2.22 5.32 3.33 7.25 8.44 5.91 6.48 4.43
TopNet [42] full 4.49 1.65 4.34 3.26 5.25 7.97 5.41 4.49 3.56
AnchorFormer [17] full 3.57 1.06 4.80 2.51 4.23 4.22 4.44 4.64 2.68
PCL2PCL [25] unpaired 16.53 4.68 28.08 6.80 17.01 20.58 14.56 28.88 11.71
KTNet [29] unpaired 6.62 1.41 5.58 4.51 8.95 8.65 8.77 9.86 5.29
LTUPCN [63] unpaired 7.35 1.67 5.78 4.21 10.35 12.55 8.84 10.49 4.96
USSPA [30] unpaired 14.08 4.01 6.94 6.81 27.36 25.39 12.13 19.31 10.70
[…]
ACL-SPC[66] no 15.26 3.38 21.72 5.87 11.38 34.73 13.03 20.24 11.73
P2C [67] no 12.28 1.85 12.73 4.93 12.99 19.43 24.82 14.95 6.57
[…]
Ours no 9.08 2.35 10.64 5.92 10.14 14.51 10.15 11.65 7.28
@@ IV-A Dataset and Evaluation Metric
To comprehensively evaluate our method, we conduct experiments using both synthetic data and real-world data. Since existing datasets lack the initial depth maps for partial point clouds, we generate a new synthetic dataset from ShapeNet [72]. Specifically, we render a single depth map for each object and then back-project it into 3D space to obtain the partial point cloud. To better align with human perspectives, we set the elevation angle for scanning shapes between -30 and 30 degrees. The complete point clouds for evaluation are uniformly sampled from the objects’ surfaces. Following PCN [15], we select 30974 objects across 8 categories: airplane, cabinet, car, chair, lamp, couch, table, and watercraft. For each category, we employ 100/150 objects for validation/testing, and the rest is used for training. We set the point numbers of partial input and complete shape to 2048 and 16384, respectively. Besides, we evaluate our model on widely used benchmarks, including PCN [15], 3D-EPN [34] and CRN [24]. The evaluation metric for synthetic data is CD. Additionally, following previous self-supervised approaches [65, 66], we report each component of CD, i.e., Precision and Coverage.
For the real-world dataset, we use the data sourced from ScanNet [73] and KITTI [74]. ScanNet contains incomplete objects in indoor scenes, and we extracted 548 chairs and 551 tables from the ScanNet dataset as indoor test data. KITTI provides partial scans of instances in outdoor scenes. We extract 473 partial point clouds from car objects in the Velodyne data, each containing at least 100 points. Due to the lack of complete ground truth, following previous works [15, 25, 27, 16, 32], we adopt the UCD and Unidirectional Hasudorff Distance (UHD) as the evaluation metrics for assessing our model’s performance on the real-world dataset. The UCD is introduced in Formula 8, and the UHD is formulated as:

### v3

@@ III Proposed Method
Denote by $\mathbf{D}_{0}\in\mathbb{R}^{H\times W}$ the raw depth map scanned from an object, $\mathbf{P}_{in}\in\mathbb{R}^{N\times 3}$ the partial point cloud back-projected by $\mathbf{D}_{0}$, $\mathbf{P}_{com}\in\mathbb{R}^{M\times 3}$ the complete shape of $\mathbf{P}_{in}$, where $H$ and $W$ represent the height and width of the depth image, $N$ and $M$ are the numbers of points. Our goal is to find a mapping to make the predicted complete shape $\mathbf{P}_{out}$ as close as possible to $\mathbf{P}_{com}$ without $\mathbf{P}_{com}$ as supervision.
As depicted in Figure 1, our framework, namely MAL-UPC, first leverages a Pattern Retrieval Network to generate complete point clouds. Then, we render the complete prediction into a silhouette map and multi-view depth maps. The rendered depth maps are used to perform adversarial learning with the single-view depth maps of the partial point clouds from the training set. In what follows, we will provide a detailed explanation of our approach.
@@ III-A Pattern Retrieval Network
The Pattern Retrieval Network leverages objects’ region-level geometric similarity to infer the missing structures. As shown in Figure 2, consuming as input a partial point cloud $\mathbf{P}_{in}$, we first use a 3D encoder derived from PCN [15] to learn a global feature $\mathbf{o}$ that encapsulates the object’s shape information. Then we employ three fully-connected layers to reconstruct a coarse yet complete point cloud $\mathbf{P}_{c}\in\mathbb{R}^{N\times 3}$. Next, we extract the position encodings and curvature encodings from the partial input and coarse output, respectively. The position encodings represent the positional relationship of local points, which are formulated as the average distance between each point and its nearest neighbors. Specifically, for each point $\mathbf{p}_{i}\in\mathbf{P}_{in}$, we first find out its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})=\{\mathbf{q}_{i}^{j}\}_{j=1}^{K}\subset\mathbf{P}_{in}$ and then compute the average Euclidean distance between $\mathbf{p}_{i}$ and $\mathcal{N}_{K}(\mathbf{p}_{i})$ to obtain its position encoding $\mathbf{f}_{pos}^{in}(i)$, i.e.,
Figure 2: The flowchart of our Pattern Retrieval Network. “FC” and “MLP” denote the Fully-Connected layer and the Multi-Layer Perceptron layer. Given a partial point cloud $\mathbf{P}_{in}$, we first utilize an auto-encoder to generate a coarse shape $\mathbf{P}_{c}$. Next, we separately extract position encodings and curvature encodings for both $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$, and then leverage a cross-attention-based module to identify and retrieve the most related encodings for each point in $\mathbf{P}_{c}$ from $\mathbf{P}_{in}$. Subsequently, $\mathbf{P}_{c}$ and the selected encodings are upsampled and fused to refine the shape.
@@
The curvature encodings capture the curvature variations across the local surfaces of the point cloud. We employ the normals of individual points to characterize surface curvature. For each point $\mathbf{p}_{i}$, we first compute the covariance matrix $\mathbf{C}_{i}$ based on its $K$ nearest neighbors $\mathcal{N}_{K}(\mathbf{p}_{i})$ determined previously. The covariance matrix captures the local surface geometry and is defined as follows:
@@
To simulate the curvature variations of the surface, inspired by [67], we calculated the cosine similarity between the normal vector of each point and its neighboring $K$ points.
@@
where $\mathbf{f}_{cur}^{in}(i)$ is the curvature encoding of $\mathbf{p}_{i}$, and $\epsilon$ is a small constant to prevent the denominator from being 0.
In this way, we obtain the position encoding $\mathbf{f}_{pos}^{in}\in\mathbb{R}^{N\times 1}$ and the curvature encoding $\mathbf{f}_{cur}^{in}\in\mathbb{R}^{N\times 1}$ from the partial input. Meanwhile, we compute the position encoding $\mathbf{f}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and the curvature encoding $\mathbf{f}_{cur}^{c}\in\mathbb{R}^{N\times 1}$ from the coarse shape using the same method. Following this, we introduce a cross-attention-based pattern retrieval module to capture correlated local patterns for coarse predictions from partial inputs. Specifically, for the position encoding, we utilize MLPs to project $\mathbf{f}_{pos}^{in}$ and $\mathbf{f}_{pos}^{c}$ to get queries $\mathbf{g}_{pos}^{c}\in\mathbb{R}^{N\times 1}$ and keys $\mathbf{h}_{pos}^{in}\in\mathbb{R}^{N\times 1}$, respectively. Then we employ the algorithm in cross-attention [51] to calculate the weight maps between $\mathbf{h}_{pos}^{in}$ and $\mathbf{g}_{pos}^{c}$, representing the correlation of the position encodings between each point in $\mathbf{P}_{in}$ and $\mathbf{P}_{c}$:
Next, based on $\mathbf{W}_{pos}$, we retrieve the $L$ most correlated position encodings from $\mathbf{f}_{pos}^{in}$ for each point in $\mathbf{P}_{c}$, denoted as $\mathbf{e}_{pos}\in\mathbb{R}^{N\times L}$. Meanwhile, the related curvature encodings $\mathbf{e}_{cur}\in\mathbb{R}^{N\times L}$ for each point in $\mathbf{P}_{c}$ are computed in the same manner. Then we upsample $\mathbf{P}_{c}$ along with $\mathbf{e}_{pos}$, $\mathbf{e}_{cur}$, $\mathbf{f}_{pos}^{c}$, $\mathbf{f}_{cur}^{c}$, $\mathbf{o}$, and fuse them in a concatenation manner. Finally, we employ MLPs to infer the point-wised offsets for the upsampled $\mathbf{P}_{c}$ and add them to obtain the final outputs $\mathbf{P}_{out}$.
@@ III-B Multi-view Adversarial Learning
Objects within the same category share many similar structures, and therefore, their shape distributions are close. To make the generated complete point cloud as realistic and reasonable as possible, we develop an adversarial learning module to learn the category-specific shape knowledge from the real depth maps.
Assuming the training set includes $U$ categories, each category comprises $G_{u}$ single-view partial point clouds to be completed and $G_{u}$ corresponding depth maps since each partial point cloud is generated by back-projecting a raw depth map into 3D space, where $u=1,2,...,U$. For each category, we gather the raw depth maps of these partial point clouds to construct a category-specific depth image bank. In other words, the depth maps in the image bank are the projections of the incomplete point clouds solely from the perspectives in which they were captured, with no additional samples included.
Next, we employ a silhouette renderer to render $\mathbf{P}_{out}$ and $\mathbf{P}_{c}$ into two silhouette maps $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ from the perspective in which $\mathbf{D}_{0}$ is captured. $\mathbf{S}_{c}$ and $\mathbf{S}_{out}$ are supervised by a mask image binarized from $\mathbf{D}_{0}$. Then we leverage a depth renderer to render $\mathbf{P}_{out}$ into depth maps from different viewpoints. These depth maps and those from the category-specific depth image bank are fed to a discriminator, which consists of five 2D convolutional layers, to discriminate whether they are generated or real samples. In practice, for each predicted result, we render a depth map from a random viewpoint at each epoch, ensuring that the viewpoints vary across all epochs to achieve multi-view rendering while saving training time.
@@
Our differentiable rendering pipeline is derived from [70]. Given the intrinsic and extrinsic camera parameters, [70] first transform the point cloud from world to camera coordinates and then project it into the 2D image plane. During rasterization, each point is rendered as a circular splat with a predefined radius in screen space, where the opacity of the splat decreases away from its center. The value of each pixel in the silhouette maps is computed by blending the information from the K-nearest points along the z-axis whose splat regions overlap that pixel. For depth maps, each pixel’s value is the smallest z-coordinate of the K-nearest points.
@@
However, [70] assigns a fixed radius to point splats under all viewpoints. In fact, the density of 2D points projected by a point cloud with a fixed point number varies across different viewpoints. For example, as shown in Figure 3, when there is a large foreground area in the rendered image, the density of the projected points is small (we call these low-density depth maps). In such cases, a small radius cannot cover all areas of the visible surface and thus many “hole pixels” appear. These “hole pixels” are inaccurately represented by the depth values of the points that should have been occluded. In contrast, in scenarios with few foreground pixels, the density of points is large (we call these high-density depth maps). In this case, a small radius suffices to cover the entire visible surface, and thus, the rendered depth map is more realistic and smooth. Therefore, a fixed point radius is inadequate for rendering under different viewpoints. Intuitively, using a larger point radius might produce smoother depth maps in both high point density and low point density scenarios. However, in this way, the projected points cover a larger area on the 2D plane, potentially extending the edges of the foreground pixels beyond the original boundaries of the object.
@@
To address this issue, we introduce a density-aware radius estimation (DARE) algorithm to adaptively determine a suitable radius for points from each viewpoint. First, we predefine an initial point radius for the reconstructed shape and then render it into a coarse depth map. We then calculate the number of foreground pixels on this depth map, denoted as $A$, and subsequently, the density of the projected points on the depth map is defined as
@@ III-C Loss Functions
Furthermore, we use a rendering loss to supervise the predicted point cloud in 2D space. First, we binarize the depth image $\mathbf{D}_{0}$ to obtain the 2D silhouette ground truth $\mathbf{S}_{0}$. Subsequently, we compute the Mean Squared Error (MSE) loss between $\mathbf{S}_{0}$ and $\mathbf{S}_{out}$. Note that due to the sparsity of the coarse point cloud $\mathbf{P}_{c}$, the foreground pixels of $\mathbf{S}_{c}$ are also sparse and discrete, whereas the foreground of $\mathbf{S}_{0}$ is continuous and dense. To solve this problem, we propose a Masked MSE loss, which only calculates the MSE loss on the foreground pixels of the rendered image. This trick confines the projected points of the rendered image to be accurately positioned.
$\begin{split}L_{rend}&=\frac{1}{H\cdot W}\sum_{i=1}^{H}\sum_{j=1}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{out}(i,j))^{2}\\ &+\frac{1}{H\cdot W}\sum_{i=1}^{H}\sum_{j=1}^{W}(\mathbf{S}_{0}(i,j)-\mathbf{S}_{c}(i,j))^{2}\cdot\mathbf{B}(i,j),\end{split}$ (10)
@@
To generate a more uniform point cloud, following [30], we incorporate a density loss $L_{dens}$ to constrain the densities of all local regions in the point cloud to be as close as possible.
@@
Methods Supervision Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
Paired Complete PC
FoldingNet [41] 6.41 1.86 9.13 3.19 5.72 12.72 5.42 7.29 6.00
PCN [15] 5.42 2.22 5.32 3.33 7.25 8.44 5.91 6.48 4.43
TopNet [42] 4.49 1.65 4.34 3.26 5.25 7.97 5.41 4.49 3.56
AnchorFormer [17] 3.57 1.06 4.80 2.51 4.23 4.22 4.44 4.64 2.68
PCL2PCL [25] 16.53 4.68 28.08 6.80 17.01 20.58 14.56 28.88 11.71
KTNet [29] 6.62 1.41 5.58 4.51 8.95 8.65 8.77 9.86 5.29
LTUPCN [63] 7.35 1.67 5.78 4.21 10.35 12.55 8.84 10.49 4.96
USSPA [30] 14.08 4.01 6.94 6.81 27.36 25.39 12.13 19.31 10.70
Ours-U 10.20 2.58 8.65 4.49 12.74 19.70 9.61 16.13 7.71
ACL-SPC [66] 15.26 3.38 21.72 5.87 11.38 34.73 13.03 20.24 11.73
P2C [67] 12.28 1.85 12.73 4.93 12.99 19.43 24.82 14.95 6.57
P2C-X 11.08 2.78 11.83 6.07 10.40 18.10 19.48 13.48 6.53
Ours-M 11.47 4.08 15.96 7.41 11.86 21.25 9.51 13.49 8.22
Ours-S 9.08 2.35 10.64 5.92 10.14 14.51 10.15 11.65 7.28
@@ IV-A Dataset and Evaluation Metric
To comprehensively evaluate our method, we conduct experiments using both synthetic data and real-world data. Since existing datasets lack the initial depth maps for partial point clouds, we generate a new synthetic dataset from ShapeNet [72]. Specifically, we render a single depth map for each object and then back-project it into 3D space to obtain the partial point cloud. The depth map size is $224\times 224$. To better align with human perspectives, we set the elevation angle for scanning shapes within the range of -30 to 30 degrees and randomly sample the azimuth angle between 0 and 360 degrees. The complete point clouds for evaluation are uniformly sampled from the objects’ surfaces. Following PCN [15], we select 30974 objects across 8 categories. For each category, we employ 100/150 objects for validation/testing, and the rest is used for training. Concretely, the object categories and their corresponding numbers of partial point clouds in the training set are as follows: airplane (3795), cabinet (1322), car (5677), chair (5750), lamp (2068), couch (2923), table (5750), watercraft (1689). Besides, we construct an image bank for each category using the raw depth maps from which the partial point clouds in the training set are derived. In total, there are 8 image banks, each containing the same number of depth maps as the partial point clouds in their corresponding category from the training set. We set the point numbers of partial inputs and complete shapes to 2048 and 16384, respectively. Besides, we evaluate our model on widely used benchmarks, including PCN [15], 3D-EPN [34] and CRN [24]. The evaluation metric for synthetic data is CD. Additionally, following previous self-supervised approaches [65, 66], we report each component of CD, i.e., Precision and Coverage.
For the real-world dataset, we use the data sourced from ScanNet [73] and KITTI [74]. ScanNet contains incomplete objects in indoor scenes, and we extracted 548 chairs and 551 tables from the ScanNet dataset as indoor test data. KITTI provides partial scans of instances in outdoor scenes. We extract 473 partial point clouds from car objects in the Velodyne data, each containing at least 100 points. Due to the lack of complete ground truth, following previous works [15, 25, 27, 16, 32], we adopt the UCD and Unidirectional Hausdorff Distance (UHD) as the evaluation metrics for assessing our model’s performance on the real-world dataset. The UCD is introduced in Formula 8, and the UHD is formulated as:
@@
IV-B Implementation Details
The point numbers of $\mathbf{P}_{in}$, $\mathbf{P}_{c}$, and $\mathbf{P}_{out}$ are 2048, 2048 and 16384, respectively. The height $H$ and width $W$ of all depth maps are 224. We employ 16 and 24 neighbors for calculating position and curvature encodings, respectively. $L$ is 16 and $\epsilon$ is $81e-$ 8. Initial point radii for the DARE algorithm are set to 0.012 for the plane category and 0.03 for other categories. The loss weights are all set to 1. The initial learning rate is 0.0001, and it decays by 0.5 for every 200 epochs. The training process takes 600 epochs to converge. The experiments are conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
IV-C Comparison with State-of-the-Art Methods
IV-C1 Results on synthetic data
We first conduct experiments on the synthetic dataset and compare the results with state-of-the-art methods, including fully supervised methods (i.e., FoldingNet [41], PCN [15], TopNet [42], AnchorFormer [17]), unpaired methods (i.e., PCL2PCL [25], KTNet [29], LTUPCN [63], USSPA [30]) and self-supervised methods (i.e., ACL-SPC [66], P2C [67]). Note that although our MAL-UPC leverages category-specific geometric knowledge to infer complete shapes and is therefore classified as an unsupervised method, we believe it is rational to compare our method against self-supervised approaches based on the following facts. First, there are no unsupervised 3D point cloud completion methods that only rely on single-view partial observations during training (i.e., our settings), except ACL-SPC, P2C, and [68]. However, the codes of [68] have not been made public yet. Regarding the weakly-supervised methods, such as [65] and [75], they use multi-view partial point clouds from complete objects for supervision. Since combining multi-view partial point clouds can nearly reconstruct a complete object [66], this supervision signal is direct and powerful, which significantly differs from our settings. Therefore, it is the most appropriate choice to compare our method with ACL-SPC and P2C. Second, it is common that common/similar geometric structures or objects exist in real-world training datasets. Thus, we propose a Multi-view Adversarial Network (MAN) to fully leverage this prior within each category, which is just our innovation. Unfortunately, ACL-SPC and P2C fail to explore this prior information. Moreover, this prior information is provided by the category-specific single-view depth maps which are 2D representations of the incomplete point clouds present in the training set. In other words, an input partial point cloud learns shape priors from depth maps of other partial point clouds to be completed in the training set, and these depth maps are from the same category as this input partial point cloud. Therefore, we do not employ any additional information beyond the training data. Finally, in addition to the self-supervised approaches, we have compared our method with other state-of-the-art unsupervised methods, i.e., unpaired methods, such as PCL2PCL, KTNet, LTUPCN and USSPA. Actually, these unsupervised methods under comparison require unpaired complete point clouds for training, and this kind of prior is stronger than that of our method.
Additionally, to provide a more comprehensive comparison, we introduce a new approach that incorporates our rendering loss and MAN into the P2C model. The weights for the rendering loss and adversarial loss are set to 1 and 0.1, respectively, while the original loss terms and weights of the P2C model remain unchanged. This experiment bridges the category-specific prior gap between the self-supervised method and our approach. Regarding our method, we present three models trained in different manners. The first model, denoted as Ours-S, involves training our MAL-UPC framework separately for each category. The second model, referred to as Ours-M, entails training MAL-UPC on all eight categories together while each input employs its category’s depth image bank for adversarial learning. The third model, Ours-U, is trained separately for each category in an unpaired manner. Specifically, the MAN module in our model is replaced with the MLP-based discriminator from USSPA, while all other components remain unchanged. For a fair comparison, the inputs to the discriminator in our unpaired model and all unpaired baselines are the predicted point cloud and a complete point cloud from our dataset that is unpaired with the partial input. Note that all compared methods are only trained separately for each category.